# Revisiting Cross-Diffusion for Overcrowding Dispersal in Interacting Species System

### v1

@@ 2 Assumptions and main results
Our first main result concerns existence of weak solution to the cross diffusion system
(2.13) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k}+\overline{f}_{k})=f_{0k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }Q,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k}+\overline{f}_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Sigma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}},\\ \\ \rho_{k}(0)=\rho_{0k}&\hbox{ in }\Omega,\end{array}\right.$
@@
  • $(\pi_{1},\pi_{2})=\pi\in H^{-1/2}(\Gamma_{N_{1}})\times H^{-1/2}(\Gamma_{N_{2}}),$
  • $(g_{1},g_{2})=g\in H^{1/2}_{00}(\Gamma_{D_{1}})\times H^{1/2}_{00}(\Gamma_{D_{2}}),$
  • for each $k=1,2,$ $(f_{0k},\overline{f}_{k})\in L^{2}(Q)\times L^{2}(Q)^{N}$,
  • $(V_{1},V_{2})=V\in L^{\infty}(Q)^{N}\times L^{\infty}(Q)^{N},$
  • $(\rho_{01},\rho_{02})=\rho_{0}\in L^{2}(Q)\times L^{2}(Q).$
@@
Under the previous assumptions, the system of PDE (2.13) has a solution $(\rho,\eta)$ in the sense that, for each $k=1,2,$ $(\rho_{k},\eta_{k})\in L^{2}_{+}(\Omega)\cap W^{1,2}(0,T;H^{-1}_{\Gamma_{D_{k}}}(\Omega)),$ $\eta_{k}\in L^{2}(0,T;H^{1}(\Omega)),$ $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$
(2.14) $\left\{\begin{array}[]{l}\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0,\end{array}\right.\quad\hbox{ a.e. in }Q,$
@@
The key idea behind the connection between $\rho$ and $\eta$ in (2.14) appears through the graph $\partial\mathcal{B}_{\alpha}(x,.),$ where for a.e. $x\in\Omega,$ $\mathcal{B}_{\alpha}(x,.)\>:\>I\!\!R^{2}\to I\!\!R^{+}$ is the application given by
$\mathcal{B}_{\alpha}(x,r)=\left\{\begin{array}[]{ll}\beta(x,\alpha_{1}\>r_{1}+\alpha_{2}\>r_{2})&\hbox{ if }r_{1},\>r_{2}\geq 0\\ \\ +\infty&\hbox{ otherwise}.\end{array}\right.$
Remember that $d\in\partial\mathcal{B}_{\alpha}(x,r)$ if and only if $d,\>r\in I\!\!R^{2}$ and $d\cdot(s-r)\leq\mathcal{B}_{\alpha}(x,s)-\beta(x,r),$ for any $s\in I\!\!R^{2}.$ Working with $\max_{m\in I\!\!R}(m\>q-\beta(x,m))=:\beta^{*}(x,q),$ the so called Legendre transform of $\beta(x,.),$ we have
@@
For a.e. $x\in\textdiv Omega,$ $\partial\mathcal{B}_{\alpha}(x,.)$ defines a maximal montone graph in $I\!\!R^{2}\times I\!\!R^{2}.$ Moreover, for any $d=(d_{1},d_{2})$ and $r=(r_{1},r_{2})\in I\!\!R^{2}$ we have
$\displaystyle d\in\partial\mathcal{B}_{\alpha}(x,r)\quad$$\Leftrightarrow$ $\displaystyle(r_{1},r_{2})\in\operatorname*{arg\,max}_{s_{1},s_{2}\geq 0}\Big{\{}s_{1}d_{1}+s_{2}d_{2}-\beta(x,\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})\Big{\}}$
@@
It is clear that $\mathcal{B}$ is convex and l.s.c, so that $\partial\mathcal{B}$ is a maximal monotone graph in $I\!\!R^{2}\times I\!\!R^{2}.$ Remember that, $d\in\partial\mathcal{B}_{\alpha}(r)$ is equivalent to say $r_{1},\>r_{2}\>\geq 0$ and
$\displaystyle r_{1}d_{1}+r_{2}d_{2}-\underbrace{\beta(x,\alpha_{1}\>r_{1}+\alpha_{2}\>r_{2})}_{\mathcal{B}_{\alpha}(x,r)}=\underbrace{\max_{s_{1},s_{2}\geq 0}\Big{\{}s_{1}d_{1}+s_{2}d_{2}-\beta(\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})\Big{\}}}_{=:\mathcal{B}_{\alpha}^{*}(x,d)},$
where $\mathcal{B}^{*}$ denotes the Fenchel conjugate of $\mathcal{B}.$ This implies the first equivalence. The remaining equivalence follows readily from the combination of two facts: the inequality
$s_{1}d_{1}+s_{2}d_{2}\leq(\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})\>\frac{d_{1}}{\alpha_{1}}\vee\frac{d_{2}}{\alpha_{2}}\leq\beta(x,\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})+\beta^{*}\left(x,\frac{d_{1}}{\alpha_{1}}\vee\frac{d_{2}}{\alpha_{2}}\right),$
holds to be true for any $d_{1},\>d_{2}\in I\!\!R$ and $s_{1},s_{2}\in I\!\!R^{+},$ and the equality ; i.e. $r_{1}d_{1}+r_{2}d_{2}=\beta(x,\alpha_{1}\>s_{1}+\alpha_{2}\>s_{2})+\beta^{*}\left(x,\frac{d_{1}}{\alpha_{1}}\vee\frac{d_{2}}{\alpha_{2}}\right),$ holds to be true for $r_{1},r_{2}\geq 0$ if and only if
(2.17) $r_{1},r_{2}\geq 0,\>r_{1}(d_{1}-d_{1}\vee d_{2})^{+}=r_{2}(d_{2}-d_{1}\vee d_{2})^{+}=0\hbox{ and }\frac{d_{1}}{\alpha_{1}}\vee\frac{d_{2}}{\alpha_{2}}\in\partial\beta\left(x,\alpha_{1}\>r_{1}+\alpha_{2}\>r_{2}\right).$
Thus the results of the proposition $\square$
@@
To illustrate the behavior of the graph $\partial\mathcal{B}_{\alpha}(x,.)$ in $I\!\!R^{2}\times I\!\!R^{2}$ and explore potential relationships between $\rho$ and $\eta,$ we present some explicit formulations of $\partial\mathcal{B}_{\alpha}(x,r)$ and its inverse $(\partial\mathcal{B}_{\alpha}(x,.))^{-1}(d),$ which coincides with $\partial\mathcal{B}_{\alpha}^{*}(x,d)$, for arbitrary $r$ and $d$ in $I\!\!R^{2}$, and a.e. $x\in\Omega.$
@@
    For a.e. $x\in\Omega,$ $\partial\mathcal{B}_{\alpha}(x,r)$ (resp. $\partial\mathcal{B}^{*}$) may be defined by
    $\partial\mathcal{B}_{\alpha}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}\in\partial\beta(0)\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(d_{1},t)\>:\>d_{1}\in\partial\beta(r_{1}),\>t\leq d_{1}\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}=0\\ \Big{\{}(t,d_{2},)\>:\>d_{2}\in\partial\beta(r_{2}),\>t\leq d_{2}\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and }r_{2}>0\\ \Big{\{}(t,t)\>:\>t\in\partial\beta(r_{1}+r_{2})\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}>0,\end{array}\right.$
    for any $r\in I\!\!R^{+}\times I\!\!R^{+},$ (resp.
    $\partial\mathcal{B}_{\alpha}^{*}(r)=\left\{\begin{array}[]{ll}\Big{\{}(0,t)\>:\>t\in\partial\beta^{*}(d_{2})\Big{\}}&\hbox{ if }d_{1}<d_{2}\\ \Big{\{}(t,0)\>:\>t\in\partial\beta^{*}(d_{1})\Big{\}}&\hbox{ if }d_{1}>d_{2}\\ \Big{\{}(t_{1},t_{2})\>:\>t_{1}\geq 0,\>t_{2}\geq 0,\>t_{1}+t_{2}\in\partial\beta^{*}(d_{1})\Big{\}}&\hbox{ if }d_{1}=d_{2}\\ \end{array}\right.$
    for any $\alpha\in I\!\!R\times I\!\!R$).
@@
    In the case where $\beta$ (resp. $\beta^{*}$) is differentiable, for a.e. $x\in\Omega,$ $\partial\mathcal{B}_{\alpha}(x,r)$ (resp. $\partial\beta^{*}(x,d)$) may be defined by
    $\partial\mathcal{B}_{\alpha}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}=0\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(\beta^{\prime}(r_{1}),t)\>:\>t\leq\beta^{\prime}(r_{1})\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}=0\\ \Big{\{}(t,\beta^{\prime}(r_{2}))\>:\>t\leq\beta^{\prime}(r_{2})\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and }r_{2}>0\\ \Big{(}\beta^{\prime}(r_{1}+r_{2}),\beta^{\prime}(r_{1}+r_{2})\Big{)}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}>0,\end{array}\right.$
    for any $r\in I\!\!R^{+}\times I\!\!R^{+}$ (resp.
    $\partial\mathcal{B}_{\alpha}^{*}(d)=\left\{\begin{array}[]{ll}(0,(\beta^{*})^{\prime}(d_{2})&\hbox{ if }d_{1}<d_{2}\\ ((\beta^{*})^{\prime}(d_{1}),0)&\hbox{ if }d_{1}>d_{2}\\ \Big{\{}(t_{1},t_{2})\>:\>t_{1}\geq 0,\>t_{2}\geq 0,\>t_{1}+t_{2}=(\beta^{*})^{\prime}(d_{1})&\hbox{ if }d_{1}=d_{2}\\ \end{array}\right.$
    for any $d\in I\!\!R\times I\!\!R$).
@@
    For instance in the case where $\beta(r)=\frac{1}{m+1}r^{m+1}$, the porous medium framework, $\partial\mathcal{B}_{\alpha}(x,r)$ may be defined by
    $\partial\mathcal{B}_{\alpha}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}=0\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(r_{1}^{m},t)\>:\>t\leq r_{1}^{m}\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and}r_{2}=0\\ \Big{\{}(t,r_{2}^{m})\>:\>t\leq r_{2}^{m}\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and}r_{2}>0\\ \Big{(}(r_{1}+r_{2})^{m},(r_{1}+r_{2})^{m}\Big{)}&\hbox{ if }r_{1}>0\hbox{ and}r_{2}>0,\end{array}\right.$
    for any $r\in I\!\!R^{+}\times I\!\!R^{+}$
@@
    In the case where $\beta(r)=I\!\!I_{[0,1]}$, which corresponds to crowed motion models, $\partial\mathcal{B}_{\alpha}(x,r)$ may be defined by
    $\partial\mathcal{B}_{\alpha}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}=0\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(\eta_{1},t)\>:\>t\leq\eta_{1}^{m},\>(1-r_{1})\>\eta_{1}=0\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}=0\\ \Big{\{}(t,\eta_{2})\>:\>t\leq\eta_{2},\>(1-r_{2})\>\eta_{2}=0\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and }r_{2}>0\\ \Big{\{}(t,t)\>:\>\>(1-r_{1}-r_{2})\>t=0\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}>0,\end{array}\right.$
    for any $r\in I\!\!R^{+}\times I\!\!R^{+}$
@@
Thanks to Remarks, the parabolic nature of the dynamics in the system (2.13) becomes clear. Indeed, assume for instance that $\overline{f}$ is equal to $0,$ and $\beta$ is differentiable. On sees that the dynamics formally split into three driver regions, each exhibiting specific behavior (assuming they are sufficiently regular):
$\left\{\begin{array}[]{lll}\rho_{1}>0,\>\rho_{2}=0\\ \\ \partial_{t}\rho_{1}-\nabla\cdot(\alpha_{1}\>\sigma_{1}\>\nabla\beta^{\prime}(\alpha_{1}\>\rho_{1})+\rho_{1}V_{1})=f_{1}\\ -\nabla\cdot(\sigma_{2}\>\nabla\eta_{2})=f_{2},\>\frac{\eta_{2}}{\alpha_{2}}\leq\frac{1}{\alpha_{1}}\beta^{\prime}(\alpha_{1}\>\rho_{1})\end{array}\right.$
$\left\{\begin{array}[]{lll}\rho_{1}=0,\>\rho_{2}>0\\ \\ -\nabla\cdot(\sigma_{1}\>\nabla\eta_{1})=f_{1},\>\frac{\eta_{1}}{\alpha_{1}}\leq\frac{1}{\alpha_{2}}\beta^{\prime}(\alpha_{2}\>\rho_{2})\\ \partial_{t}\rho_{2}-\nabla\cdot(\alpha_{2}\>\sigma_{2}\>\nabla\beta^{\prime}(\alpha_{2}\>\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
$\left\{\begin{array}[]{lll}\partial_{t}(\rho_{1}-\rho_{2})+\nabla\cdot(\rho_{1}\>V_{1}-\rho_{2}\>V_{2})=f_{1}-f_{2}\\ \\ \partial_{t}\rho_{1}-\nabla\cdot(\alpha_{1}\>\sigma_{1}\>\nabla\beta^{\prime}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})+\rho_{1}V_{1})=f_{1}\\ \partial_{t}\rho_{2}-\nabla\cdot(\alpha_{2}\>\sigma_{2}\>\nabla\beta^{\prime}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
@@
(2.18) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{1}-\nabla\cdot(\alpha_{1}\>\sigma_{1}\>\rho_{1}\>\nabla\varphi^{\prime}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})+\rho_{1}\>V_{1})=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\alpha_{2}\>\sigma_{2}\>\rho_{2}\>\nabla\varphi^{\prime}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})+\rho_{1}\>V_{1})=f_{2}.\end{array}\right.$
Both systems share a similar behavior in the segregated regions, $S_{1},$ and $S_{2}$ as long as the function $\varphi$ satisfies $\beta(r)=r\>\varphi(r)-r,$ for any $r\in I\!\!R.$ However, there are key differences. For instance in the case where $\alpha_{k}=1,$ for $k=1,2,$ :

### v2

@@ 2 Assumptions and main results
    When dealing with the case where $\upsilon\in L^{p^{\prime}}(\Omega)$ and $-\nabla\cdot\upsilon=f\in H^{-1}_{\tilde{\Gamma}}(\Omega),$ defining the trace of $\upsilon$ on the boundary can be tricky. In general, it might not have a well-defined meaning. Around this, the trace $(\upsilon+\overline{f})\cdot\nu$ is well defined on $\partial\Omega,$ and does not depend on the specific choice $\overline{f}$ in the decomposition $f=f_{0}+\nabla\cdot\overline{f}$ given by (2.12). Therefore, when working with mixed boundary conditions like equation (2.11), it becomes crucial to consider the individual boundary trace of each $f_{k}$ on $\Gamma_{N_{k}}.$
2.1 Existence result
To treat (2.10) for general source terms $f_{k}$ in the dual space $H^{-1}_{\Gamma_{D_{k}}}(\Omega)$, $k=1,2,$ we consider the following cross-diffusion system
(2.15) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k}+\overline{f}_{k})=f_{0k},\\ \\ \eta_{k}\in\partial\beta(x,\rho_{1}+\rho_{2})+\partial I\!\!I_{[0,\infty)}(\rho_{k})\end{array}\right\}&\hbox{ in }Q,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k}+\overline{f}_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Sigma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}},\\ \\ \rho_{k}(0)=\rho_{0k}&\hbox{ in }\Omega,\end{array}\right\}\hbox{ for }k=1,2$
@@ 2.1 Existence result
  • $(\pi_{1},\pi_{2})=\pi\in H^{-1/2}(\Gamma_{N_{1}})\times H^{-1/2}(\Gamma_{N_{2}})$
  • $(g_{1},g_{2})=g\in H^{1/2}_{00}(\Gamma_{D_{1}})\times H^{1/2}_{00}(\Gamma_{D_{2}})$
  • $(f_{1},f_{2})=f\in L^{2}(0,T;H^{-1}_{\Gamma_{D_{1}}}(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}(\Omega))$, where each $f_{k}$ is identified with the couple $(f_{0k},\overline{f}_{k})\in L^{2}(Q)\times L^{2}(Q)^{N},$ for each $k=1,2,$ as given by (2.12).
  • $(V_{1},V_{2})=V\in L^{\infty}(Q)^{N}\times L^{\infty}(Q)^{N}$
  • $(\rho_{01},\rho_{02})=\rho_{0}\in L^{2}(Q)\times L^{2}(Q),$ assumed to be such that
    (2.16) $\int_{\Omega}\beta(.,\rho_{01}+\rho_{02})<\infty.$
Our first main result concerns existence of weak solution to (2.15).
@@
Under the previous assumptions, the system of PDE (2.15) has a solution $(\rho,\eta)$ in the sense that, for each $k=1,2,$ $\rho_{k}\in L^{2}_{+}(Q)\cap W^{1,2}(0,T;H^{-1}_{\Gamma_{D_{k}}}(\Omega)),$ $\rho_{k}(0)=\rho_{0k},$ $\eta_{k}\in L^{2}(0,T;H^{1}(\Omega)),$ $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$
(2.17) $\eta_{1}\vee\eta_{2}=:\tilde{\eta}\in\partial\beta(x,\rho_{1}+\rho_{2}),\>\quad\eta_{k}-\tilde{\eta}\in\partial I\!\!I_{[0,\infty)}(\rho_{k}),\quad\hbox{ a.e. in }Q,$
and
@@
While a detailed improvement will be provided in Section 3, we observe here that the connection between $\rho$ and $\eta$ in (2.17) hinges essentially on the graph $\partial\mathcal{B}(x,.),$ where for a.e. $x\in\Omega,$ $\mathcal{B}(x,.)\>:\>I\!\!R^{2}\to I\!\!R^{+}$ is the application given by
$\mathcal{B}(x,r)=\left\{\begin{array}[]{ll}\beta(x,r_{1}+r_{2})&\hbox{ if }r_{1},\>r_{2}\geq 0\\ \\ +\infty&\hbox{ otherwise}.\end{array}\right.$
Remember that $d\in\partial\mathcal{B}(x,r)$ if and only if $d,\>r\in I\!\!R^{2}$ and $d\cdot(s-r)\leq\mathcal{B}(x,s)-\mathcal{B}(x,r),$ for any $s\in I\!\!R^{2}.$ Working with $\max_{m\in I\!\!R^{2}}(m\cdot q-\beta(x,m))=:\beta^{*}(x,q),$ the so called Legendre transform of $\beta(x,.),$ we have
@@
For a.e. $x\in\Omega,$ $\partial\mathcal{B}(x,.)$ defines a maximal montone graph in $I\!\!R^{2}\times I\!\!R^{2}.$ Moreover, for any $d=(d_{1},d_{2})$ and $r=(r_{1},r_{2})\in I\!\!R^{2}$ we have
$\displaystyle d\in\partial\mathcal{B}(x,r)\quad$$\Leftrightarrow$ $\displaystyle(r_{1},r_{2})\in\operatorname*{arg\,max}_{s_{1},s_{2}\geq 0}\Big{\{}s_{1}d_{1}+s_{2}d_{2}-\beta(x,s_{1}+s_{2})\Big{\}}$
[…]
@@
It is clear that, for a.e. $x\in\Omega,$ $\mathcal{B}(x,.)$ is convex and l.s.c, so that $\partial\mathcal{B}(x,.)$ is a maximal monotone graph in $I\!\!R^{2}\times I\!\!R^{2}.$ Remember that, $d\in\partial\mathcal{B}(x,r)$ is equivalent to say $r_{1},\>r_{2}\>\geq 0$ and
$\displaystyle r_{1}d_{1}+r_{2}d_{2}-\underbrace{\beta(x,r_{1}+r_{2})}_{\mathcal{B}(x,r)}=\underbrace{\max_{s_{1},s_{2}\geq 0}\Big{\{}s_{1}d_{1}+s_{2}d_{2}-\beta(s_{1}+s_{2})\Big{\}}}_{=:\mathcal{B}^{*}(x,d)},$
where $\mathcal{B}^{*}(x,.)$ denotes the Fenchel conjugate of $\mathcal{B}(x,.).$ This implies the first two equivalence assertions. The remaining parts follows readily from the combination of two facts: the inequality
$s_{1}d_{1}+s_{2}d_{2}\leq(s_{1}+s_{2})\>d_{1}\vee d_{2}\leq\beta(x,s_{1}+s_{2})+\beta^{*}\left(x,d_{1}\vee d_{2}\right),$
holds to be true for any $d_{1},\>d_{2}\in I\!\!R$ and $s_{1},s_{2}\in I\!\!R^{+},$ and the equality ; i.e. $r_{1}d_{1}+r_{2}d_{2}=\beta(x,s_{1}+s_{2})+\beta^{*}\left(x,d_{1}\vee d_{2}\right),$ holds to be true for $r_{1},r_{2}\geq 0$ if and only if
(2.20) $r_{1},r_{2}\geq 0,\>r_{1}\left(d_{2}-d_{1}\vee d_{2}\right)^{+}=r_{2}\left(d_{1}-d_{1}\vee d_{2}\right)^{+}=0\hbox{ and }d_{1}\vee d_{2}\in\partial\beta\left(x,r_{1}+r_{2}\right).$
$\square$
@@
To illustrate the relationships between the densities $\rho_{1},\>\rho_{2},$ and the potentials $\eta_{1},\>\eta_{2},$ let us explore potential behaviors of the graph $\partial\mathcal{B}(x,.)$ in $I\!\!R^{2}\times I\!\!R^{2}.$ We present some explicit formulations of $\partial\mathcal{B}$ and its inverse $(\partial\mathcal{B}(x,.))^{-1},$ which coincides with $\partial\mathcal{B}^{*}(x,.).$
@@
    For a.e. $x\in\Omega,$ (.$\partial\mathcal{B}(.,r)$
    $\partial\mathcal{B}(x,r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}\in\partial\beta(x,0)\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(d_{1},t)\>:\>d_{1}\in\partial\beta_{1}(x,r_{1}),\>t\leq d_{1}\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}=0\\ \Big{\{}(t,d_{2})\>:\>d_{2}\in\partial\beta(x,r_{2}),\>t\leq d_{2}\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and }r_{2}>0\\ \Big{\{}(t,t)\>:\>t\in\partial\beta(x,r_{1}+r_{2})\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}>0,\end{array}\right.\quad\forall r\in I\!\!R^{+}\times I\!\!R^{+}$
    and
    $\partial\mathcal{B}^{*}(x,d)=\left\{\begin{array}[]{ll}\Big{\{}(0,t)\>:\>t\in\partial\beta^{*}(x,d_{2})\Big{\}}&\hbox{ if }d_{1}<d_{2}\\ \Big{\{}(t,0)\>:\>t\in\partial\beta^{*}(x,d_{1})\Big{\}}&\hbox{ if }d_{1}>d_{2}\\ \Big{\{}(t_{1},t_{2})\>:\>t_{1}\geq 0,\>t_{2}\geq 0,\>t_{1}+t_{2}\in\partial\beta^{*}(d_{1})\Big{\}}&\hbox{ if }d_{1}=d_{2}\\ \end{array}\right.\quad\forall d\in I\!\!R\times I\!\!R.$
@@
    In the case where $\beta(x,.)$ (resp. $\beta_{1}^{*}(x,.)$) is differentiable, for a.e. $x\in\Omega,$ we have
    $\partial\mathcal{B}(x,r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}=0\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(\beta^{\prime}(x,r_{1}),t)\>:\>t\leq\beta^{\prime}(x,r_{1})\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}=0\\ \Big{\{}(t,\beta^{\prime}(x,r_{2}))\>:\>t\leq\beta^{\prime}(x,r_{2})\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and }r_{2}>0\\ \Big{(}\beta^{\prime}(r_{1}+r_{2}),\beta^{\prime}(x,r_{1}+r_{2})\Big{)}&\hbox{if }r_{1}>0\hbox{ and }r_{2}>0,\end{array}\right.\quad\forall r\in I\!\!R^{+}\times I\!\!R^{+}$
    and
    $\partial\mathcal{B}^{*}(x,d)=\left\{\begin{array}[]{ll}(0,(\beta^{*})^{\prime}(x,d_{2})&\hbox{ if }d_{1}<d_{2}\\ ((\beta^{*})^{\prime}(x,d_{1}),0)&\hbox{ if }d_{1}>d_{2}\\ \Big{\{}(t_{1},t_{2})\>:\>t_{1}\geq 0,\>t_{2}\geq 0,\>t_{1}+t_{2}=(\beta^{*})^{\prime}(x,d_{1})&\hbox{ if }d_{1}=d_{2}\\ \end{array}\right.\quad\forall d\in I\!\!R\times I\!\!R.$
@@
    For instance in the case where $\beta(r)=\frac{1}{m+1}r^{m+1}$ (porous medium equation) $\partial\mathcal{B}(r)$ may be defined by
    $\partial\mathcal{B}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}=0\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(r_{1}^{m},t)\>:\>t\leq r_{1}^{m}\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and}r_{2}=0\\ \Big{\{}(t,r_{2}^{m})\>:\>t\leq r_{2}^{m}\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and}r_{2}>0\\ \Big{(}(r_{1}+r_{2})^{m},(r_{1}+r_{2})^{m}\Big{)}&\hbox{ if }r_{1}>0\hbox{ and}r_{2}>0,\end{array}\right.\quad\forall r\in I\!\!R^{+}\times I\!\!R^{+}.$
@@
    In the case where $\beta=I\!\!I_{[0,1]}$, which corresponds to crowed motion models, $\partial\mathcal{B}(r)$ may be defined by
    $\partial\mathcal{B}(r)=\left\{\begin{array}[]{ll}\Big{\{}(d_{1},d_{2})\>:\>d_{1}\vee d_{2}=0\Big{\}}&\hbox{ if }r_{1}=r_{2}=0\\ \Big{\{}(\eta_{1},t)\>:\>t\leq\eta_{1},\>\eta_{1}\in\partial I\!\!I_{[0,1]}(r_{1})\Big{\}}&\hbox{ if }r_{1}>0\hbox{ and }r_{2}=0\\ \Big{\{}(t,\eta_{2})\>:\>t\leq\eta_{2},\>\eta_{2}\in\partial I\!\!I_{[0,1]}(r_{2})\Big{\}}&\hbox{ if }r_{1}=0\hbox{ and }r_{2}>0\\ \Big{\{}(t,t)\>:\>\>t\in\partial I\!\!I_{[0,1]}(r_{1}+r_{2})\Big{\}}&\hbox{ if}r_{1}>0\hbox{ and }r_{2}>0,\end{array}\right.\quad\forall r\in I\!\!R^{+}\times I\!\!R^{+}.$
@@
Assume for instance that $\overline{f}$ is equal to $0,$ and $\beta$ is differentiable. On sees that the dynamic of the system (2.15) splits formally into three driver regions, $S_{1}:=[\rho_{1}>0\>\>\&\>\>\rho_{2}=0]$, $S_{2}:=[\rho_{1}=0\>\>\&\>\>\rho_{2}>0]$ and $S:=[\rho_{1}>0\>\>\&\>\>\rho_{2}>0]$ each exhibiting specific behavior. Assuming they are sufficiently regular, we can formally observe the following behavior
${S_{1}=[\rho_{1}>0\>\&\>\rho_{2}=0]}$$\left\{\begin{array}[]{lll}\partial_{t}\rho_{1}-\nabla\cdot(\sigma_{1}\>\nabla\beta^{\prime}(\rho_{1})+\rho_{1}V_{1})=f_{1}\\ \\ -\nabla\cdot(\sigma_{2}\>\nabla\eta_{2})=f_{2},\>\eta_{2}\leq\beta^{\prime}(\rho_{1})\end{array}\right.$
${S_{2}=[\rho_{1}=0\>\&\>\rho_{2}>0]}$$\left\{\begin{array}[]{lll}-\nabla\cdot(\sigma_{1}\>\nabla\eta_{1})=f_{1},\>\eta_{1}\leq\beta^{\prime}(\rho_{2})\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\sigma_{2}\>\nabla\beta^{\prime}(\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
${S=[\rho_{1}>0\>\&\>\rho_{2}>0]}$$\left\{\begin{array}[]{lll}\partial_{t}\rho_{1}-\nabla\cdot(\sigma_{1}\>\nabla\beta^{\prime}(\rho_{1}+\rho_{2})+\rho_{1}V_{1})=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\sigma_{2}\>\nabla\beta^{\prime}(\rho_{1}+\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
@@
(2.21) $\left\{\begin{array}[]{ll}\partial_{t}\rho_{1}-\nabla\cdot(\sigma_{1}\>\rho_{1}\>\nabla\varphi^{\prime}(\rho_{1}+\rho_{2})+\rho_{1}\>V_{1})=f_{1}\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\sigma_{2}\>\rho_{2}\>\nabla\varphi^{\prime}(\rho_{1}+\rho_{2})+\rho_{1}\>V_{1})=f_{2}.\end{array}\right.$
Both systems share some similar behaviors in the segregated regions, $S_{1},$ and $S_{2}$ as long as the function $\varphi$ satisfies $\beta(r)=r\>\varphi(r)-r,$ for any $r\in I\!\!R.$ However, there are key differences. For instance :